A rectangular room is 1.4 times as long as it is wide, and its perimeter is 27 meters. Find the dimension of the room.
Answer by macston(5194)

W=width; L=length=1.4(W); P=perimeter=2(L+W)
P=2(L+W)=27 Substitute for L
2(1.4(W)+W)=27 meters
2.8W+2W=27 meters
4.8W=27 meters Divide each side by 4.8
W=5.625 meters ANSWER 1: Width is 5.625 meters
L=1.4(W)=1.4(5.625 meters)=7.875 meters ANSWER 2: Length is 7.875 meters
CHECK
P=2(L+W)=27 meters
2(7.875 meters + 5.625 meters)=27 meters
2(13.5 meters=27 meters
27 meters=27 meters